OpenStudy (anonymous):

Use substitution to find the solutions (if any). 3x + 7y = 3 and 12x + 28y = 16

OpenStudy (anonymous):

u know how 2 do substitution?

OpenStudy (saifoo.khan):

they are parallel lines, they wont meet. so there wont be any solution

OpenStudy (anonymous):

still learning

OpenStudy (vijay):

many solutions

OpenStudy (anonymous):

i need to show my work

OpenStudy (anonymous):

Okay- the goal of substitution is to get one equation with one variable. First, you put one of the equations in terms of one of the variables. i.e, if your equation was a+b=7, and u wanted 2 put it in terms of a, it would become a=7-b

OpenStudy (saifoo.khan):

vijay, u r incorrect.

OpenStudy (anonymous):

Then, you'd put 7-b everywhere you saw a in the second equation. That's substitution

OpenStudy (anonymous):

SO for this problem, get an equation in terms of y or x. lets do x, caus x's are cool.

OpenStudy (anonymous):

this is the answer I got for the first 1 x = 1 + -2.333333333y

OpenStudy (anonymous):

3x+7y=3 3x=3-7y x=1-(7/3y)

OpenStudy (anonymous):

I'd leave it in fractional form, caus it looks a lot nicer that way.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so plug 1-(7/3 y) wherever u c x in the second equation I'd use parenthesis in case u have 2 distribute (1-(7-3 y))

OpenStudy (anonymous):

sry! meant 7/3!

OpenStudy (anonymous):

3((1-(7/3y))+2y/3=8 right?

OpenStudy (anonymous):

hold on-

OpenStudy (anonymous):

Not quite sure what u did. u started with 12(1-(7/3 y))+28y=16, right?

OpenStudy (anonymous):

sorry wrong prob...lol

OpenStudy (anonymous):

for me not you

OpenStudy (anonymous):

no hugie!

OpenStudy (anonymous):

i have been doing this for 10 hours I am brain dead

OpenStudy (anonymous):

THAT stinks :(. Pull through and I'll give u a medal!

OpenStudy (anonymous):

this is where I have the problem is on the second equation

OpenStudy (anonymous):

ok So we have 12(1-(7/3 y))+28y=16,right? I'd divide the whole equation by 4 to simplify it. just means u have 3(1-(7/3 y))+7y=4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

And then distribute the negative sign 3(1(-7/3 y))+7y=4

OpenStudy (anonymous):

distribute the 1 (doesn't change anything), and then the 3. -7y+7y=4

OpenStudy (anonymous):

you lost me there

OpenStudy (anonymous):

which step?

OpenStudy (anonymous):

distribute one

OpenStudy (anonymous):

okay. u know how 2 distribute, right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

show me what you mean..can yu write it out

OpenStudy (anonymous):

Okay then. THere's something called Distributive Property in math. It just means multiplying a number through a pair of parenthesis. 3(x-1) Distribute the 3 3x-3 multiply the 3 by everything in the parenthesis 3*x=3x 3*-1=-3 3x-3

OpenStudy (anonymous):

geddit or want me to show u another one?

OpenStudy (anonymous):

no i think i got it

OpenStudy (anonymous):

so my answer is 4

OpenStudy (anonymous):

no. 3(1-(7/3 y)+7y=4 first distribute the negative sign through cool math fact- negative signs can be rewritten as -1 times the number being minus-ed -3=-1*3 that means multiply every term in the parenthesis by a negative one

OpenStudy (anonymous):

so i get 3x.666666667=7y=4?

OpenStudy (anonymous):

i ment +

OpenStudy (anonymous):

Not done yet 3(1(-7/3 y) +7y=4 Now distribute the 1 but anything multiplied by a positive one stays the same number ex. 6*1=6 so now we have 3(-7/3 y)+7y=4

OpenStudy (anonymous):

Distribute the 3, and your denominator is no more -7*3 ----- 3's cancel out. 3

OpenStudy (anonymous):

Now u have -7y+7y=4

OpenStudy (anonymous):

so 3*-7+7y=4?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so now we have y=4

OpenStudy (anonymous):

No, not quite. now we have -7y+7y=4 7y-7y=4 0=4 it doesn't work- there's no solution All that work for nothing! (literally)

OpenStudy (anonymous):

ok now I have to A. For equation ___________ 1. The x-intercept is ___. 2. The y-intercept is ___. 3. The slope is ___. B. For equation ___________ 1. The x-intercept is ___. 2. The y-intercept is ___. 3. The slope is ___.

OpenStudy (anonymous):

can you help with this?

OpenStudy (anonymous):

Get each equation into y=mx+b form u know what that is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. Put each equation into that form and post them

OpenStudy (anonymous):

i am so tired I cant figure it out. I will just stop here and turn this in like it is.

OpenStudy (anonymous):

COme on, ur so close!

OpenStudy (anonymous):

do the first one...that will be a hint for me

OpenStudy (anonymous):

Hm- How about I do the second and u do the first? (leave the slope in fractional form)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

second equation y=-3/7 x +2/7 what's the first?

OpenStudy (anonymous):

y= i dont know

OpenStudy (anonymous):

first equation y=-3/7x+3

OpenStudy (anonymous):

First equation slope: -3/7 x y intercept: 3 x intercept:7

OpenStudy (anonymous):

you have no idea how much I appreciate all your help

OpenStudy (anonymous):

i got 11 out of 55 on my last paper

OpenStudy (anonymous):

Second equation slope: -3/7 y intercept: 2/7 x intercept:-2/3 drop the x for the first slope

OpenStudy (anonymous):

good luck. Only gave u the answers b/c u seemed desperate

OpenStudy (anonymous):

I have 3 word problems left then I can quit

OpenStudy (anonymous):

trust me at this point i am

OpenStudy (anonymous):

thanks so much!!!!!!! for all your help

OpenStudy (anonymous):

welcome g ot 2 get off